### v1

# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, for example, in industrial contexts.
1 Introduction
Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$. Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ (Hawkins, 1980)—often indicating critical, rare, or unforeseen events. The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security (Siddiqui et al., 2019), healthcare (Fernando et al., 2021; Tian et al., 2023), and industrial inspection. In this work, we focus on the latter, where fully automated systems necessitate the ability to detect defective or missing parts to prevent malfunctions in downstream products, raise alerts for potential hazards, or analyze these to optimize production lines. See the right-hand side of Figure 1 for anomalous samples in this context (with predicted anomaly maps from AnomalyDINO).
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). Whenever suitable we select the relevant patches from the (potentially augmented) reference sample via masking and collect the nominal patch representations in $\mathcal{M}$. For both, masking and feature extraction, we utilize DINOv2. At test time, the distances of patch representations to those in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score after aggregation with $q$. Further examples for other categories are depicted on the right (and in Figures 3 and 4 in Appendix A).
AD for industrial images has gained tremendous interest over the last couple of years. The close-to-optimal results on benchmark data make it seem that the problem of anomaly detection is essentially solved. For instance, Mousakhan et al. (2023) report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD (Bergmann et al., 2021) and VisA (Zou et al., 2022), respectively. The most popular AD techniques use the training data to train an anomaly classifier (e.g., Ruff et al., 2020), or a generative model coupled with reconstruction-based (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Mousakhan et al., 2023) or likelihood-based (e.g., Rudolph et al., 2022; Daehyun Kim, 2023) anomaly scoring. However, these approaches operate within the full-shot setting, meaning they rely on access to a sufficiently large amount of training data. Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution (Li et al., 2024a), there is an increasing interest in few-shot and zero-shot anomaly detection.
Few-shot techniques, however, heavily rely on meaningful features, or as Reiss et al. (2022) frame it: ‘anomaly detection requires better representations’. Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion (e.g., Radford et al., 2021; Caron et al., 2021; Oquab et al., 2024). The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches (e.g., Jeong et al., 2023; Cao et al., 2023; Zhou et al., 2024; Kwak et al., 2024)
Here, we propose to focus on a vision-only approach, in contrast to such multimodal techniques. This perspective is motivated by the observation that few-shot anomaly detection is feasible for human annotators based on visual features only, and does not require additional textual description of the given object or the expected types of anomalies (which are typically not known a priori).
Our approach, termed AnomalyDINO, is memory-bank-based and leverages DINOv2 (Oquab et al., 2024) as a backbone (but any architecture with strong patch-level feature extraction is suitable for the proposed approach). We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by Chen et al. (2023) or Kwak et al. (2024). Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD (Bergmann et al., 2021) and outperforms all but one competing method on VisA (Zou et al., 2022).
The structure of the paper is as follows: Section 2 reviews relevant prior studies and clarifies the distinctions between the settings addressed by zero- and few-shot, and batched zero-shot techniques. Section 3 introduces our proposed method, AnomalyDINO. An extension of this method to the batched zero-shot scenario is detailed in Appendix D. Section 4 presents the experimental outcomes. Additional results and an ablation study are provided in Appendices A and C, respectively. We address identified failure cases of AnomalyDINO in Appendix B.
Contributions
  • We propose AnomalyDINO, a simple and training-free yet highly effective patch-based technique for visual anomaly detection. Our method builds upon the high-quality feature representation extracted by DINOv2.
  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques, thereby reducing the gap between few- and full-shot anomaly detection. Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6%. Moreover, our results on VisA are not only competitive with other few-shot methods but also establish new state-of-the-art performance among all training-free few-shot anomaly detectors.
2 Related Work
Foundation Models for Vision
Multimodal foundation models have emerged as powerful tools for a wide range of tasks (see e.g., Bommasani et al., 2022; Li et al., 2023; Chen et al., 2020; He et al., 2022; Radford et al., 2021; Liu et al., 2023; OpenAI, 2024). Most relevant to visual AD are multimodal approaches based on CLIP (Radford et al., 2021) or recent LLMs (e.g., OpenAI, 2024), but also vision-only approaches like DINO (Caron et al., 2021; Oquab et al., 2024). CLIP (Radford et al., 2021) learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO (Caron et al., 2021; Oquab et al., 2024) leverages a self-supervised student-teacher framework based on vision transformers (Dosovitskiy et al., 2021). It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 (Oquab et al., 2024) combines ideas from DINO with patch-level reconstruction techniques (Zhou et al., 2022) and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO (Liu et al., 2023), builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
Anomaly Detection
Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept (Yang et al., 2021; Ruff et al., 2021). In this work, we focus on low-level sensory anomalies of industrial image data, i.e., we do not target the detection of semantic anomalies but of low-level features such as scratches of images of industrial products (see e.g., Figure 1). Several works tackled this task by either training an anomaly classifier (e.g., Ruff et al., 2020) or a generative model, which allows for reconstruction-based or likelihood-based AD (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Zhang et al., 2023; Daehyun Kim, 2023; Mousakhan et al., 2023).
However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples. Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD. The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets (He et al., 2015), Wide ResNets (Zagoruyko and Komodakis, 2016), and Vision Transformers (ViT, Dosovitskiy et al., 2020). These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set. The comparison with the memory bank is typically conducted using a nearest neighbor approach (Bergman et al., 2020; Cohen and Hoshen, 2020; Defard et al., 2020; Roth et al., 2022; Xie et al., 2023; Li et al., 2024b). Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings (Jeong et al., 2023; Chen et al., 2023; Kwak et al., 2024; Zhou et al., 2024). Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., Chen et al., 2023, use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs (Xu et al., 2024). Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
Categorization of Few-/Zero-Shot Anomaly Detectors
Previous works consider different AD setups, which complicates their evaluation and comparison. To remedy this, we provide a taxonomy of recent few- and zero-shot AD based on the particular ‘shot’-setting, the training requirements, and the modes covered by the underlying models. We categorize three ‘shot’-settings: zero-shot, few-shot, and batched zero-shot. Zero- and few-shot settings are characterized by the number of nominal training samples a method can process, before making predictions on the test samples. In batched zero-shot, inference is not performed sample-wise but based on a whole batch of test samples, usually the full test set. For instance, the method proposed by Li et al. (2024b) benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images. Another work that considers this setting is that of Li et al. (2024a), who deploy a parameter-free anomaly detector based on the effect of batch normalization. We split the training requirements into the categories ‘Training-Free’, ‘Fine-Tuning’, and ‘Meta-Training’. ‘Training-Free’ approaches do not require any training, while ‘Fine-Tuning’ methods use the few accessible samples to modify the underlying model. In contrast, ‘Meta-Training’ is associated with training the model on a dataset related to the test data. For example, Li et al. (2024a) train their model on MVTec-AD containing all classes except the class they test against. Zhou et al. (2024) and Chen et al. (2023) train their model on VisA when evaluating the test performance on MVTec-AD and vice versa. Finally, we differentiate the leveraged models, which are either vision models (such as pre-trained ViT) or language-vision models (such as CLIP). We provide a detailed summary in Table 1.
Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods. The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., Roth et al. (2022).
Method Setting Training Type Modes
DN2 (Bergman et al., 2020) Few-Shot Training-Free Vision
SPADE (Cohen and Hoshen, 2020) Few-Shot Training-Free Vision
PaDiM (Defard et al., 2020) Few-Shot Fine-Tuning Vision
PatchCore (Roth et al., 2022) Few-Shot Training-Free Vision
PatchCore–opt (Santos et al., 2023) Few-Shot Training-Free Vision
MuSc (Li et al., 2024b) Batched Zero-Shot Training-Free Vision
GraphCore (Xie et al., 2023) Few-Shot Training (GNN) Vision
WinCLIP (Jeong et al., 2023) Zero-/Few-Shot Training-Free Vision + Language
APRIL-GAN (Chen et al., 2023) Zero-/Few-Shot Meta-Training Vision + Language
ADP (Kwak et al., 2024) Few-Shot Fine-Tuning Vision + Language
AnomalyCLIP (Zhou et al., 2024) Zero-Shot Meta-Training Vision + Language
GPT4-V (Xu et al., 2024) Zero-Shot Training-Free Vision + Language
ACR (Li et al., 2024a) Batched Zero-Shot Meta-Training /
AnomalyDINO (Ours) Few-Shot Training-Free Vision
3 Patch-Matching for Visual Anomaly Detection
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data. We follow a memory-bank approach, i.e., we first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
3.1 Anomaly Detection with Nearest Patch Distances
Let us consider the image space $\mathcal{X}$. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{F}$ denotes the feature space. Note that $n$ depends on the image resolution and the patch size. Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in\{1,\dots,k\}\bigr{\}}$ ($k$-shot setting), we collect the nominal patch features and store them in a memory bank
$\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in\{1,\dots,n\}\bigr{\}}\enspace.$ (1)
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with our memory bank $\mathcal{M}$. To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
$d_{\operatorname{NN}}(\mathbf{p};\mathcal{M}):=\min_{\mathbf{p}_{\operatorname{ref}}\in\mathcal{M}}d(\mathbf{p},\,\mathbf{p}_{\operatorname{ref}})$ (2)
for some distance metric $d$. In our experiments, we set $d$ as the cosine distance, that is,
$d(\mathbf{x},\mathbf{y}):=1-\frac{\langle\mathbf{x},\;\mathbf{y}\rangle}{\|\mathbf{x}\|\,\|\mathbf{y}\|}\enspace.$ (3)
The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is given by aggregating the patch distances via a suitable statistic $q$
$s(\mathbf{x}_{\mathrm{test}}):=q\big{(}\{d_{\operatorname{NN}}(\mathbf{p}_{1};\mathcal{M}),\dots,d_{\operatorname{NN}}(\mathbf{p}_{n};\mathcal{M})\}\big{)}\enspace.$ (4)
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile (e.g., McNeil, 1999) and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation.
Following Roth et al. (2022), we utilize bilinear upsampling and Gaussian smoothing (with $\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 3 and 4).
We extend AnomalyDINO also to the batched zero-shot scenario, see Appendix D (Figure 13).
3.2 Enriching the Memory Bank & Filtering Relevant Patches
Access to only a single/few normal sample(s) poses (at least) two challenges. First, all relevant information about the concept of normality must be extracted from the single/few available nominal sample(s). However, it is expected that the variability at test time is (considerably) higher than in $X_{\mathrm{ref}}$. Secondly, one must determine a suitable preprocessing for reference and test samples, solely based on $X_{\mathrm{ref}}$. To address low variations of nominal patch features, we apply simple augmentations in the form of rotations, following the insights of Xie et al. (2023). In addition, to ensure that $\mathcal{M}$ is of high quality, we consider masking the object of interest as background noise could poison $\mathcal{M}$. Note, that a suitable preprocessing, i.e., whether to apply masking or rotations, should depend on the object of interest (see, e.g., Figures 7 and 10, and Section C.1 for further discussion).
Figure 2: Masking test on MVTec-AD. The dashed red regions mark areas that belong to the object of interest but are falsely predicted to be background. See App. C.1 for the outcomes per category.
Masking
To minimize the overhead of the proposed method, we utilize DINOv2 also for masking, i.e., delineating the primary object(s) in an image from its background. This is achieved by thresholding the first PCA component of the patch features (Oquab et al., 2024). We observe empirically that this procedure sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches, see Figure 2. To resolve that issue we test whether the PCA-based mask correctly captures the object in the first reference sample and apply masking accordingly. In general, we do not mask textures (e.g., wood or tile in MVTec-AD).
Rotation
Rotating the reference sample may improve the detection performance by enriching the memory bank, but can also be detrimental in cases where rotations of (parts of the) object of interest can be considered anomalies themselves. We consider two different settings.
In the ‘agnostic’ case, which we focus on in the main paper, we always augment the reference sample with rotations. We also consider the case where we know about the samples’ potential rotations (‘informed’). The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to significantly lower inference time. See Section C.1 for the effect of masking and rotation, and a comparison of ‘informed’ vs. ‘agnostic’.
4 Experiments
AnomalyDINO – Defaults
Using DINOv2111Code and model weights of DINOv2 (Oquab et al., 2024) are available under Apache 2.0 license. as backbone allows us to choose from the distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters). To prioritize low latency, we take the smallest model (S) as our default. As discussed above, the ‘agnostic’ preprocessing is our default.
Datasets
For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset (Bergmann et al., 2021) is available under the CC BY-NC-SA 4.0 license. and VisA, [^3] The VisA dataset (Zou et al., 2022) is released under the CC BY 4.0 license. two datasets with high-resolution images for industrial anomaly detection. MVTec-AD consists of fifteen categories, both objects and textures, with a small amount of clean training samples (between 60 and 391), and up to eight anomaly types per category. VisA has twelve categories with up to 1006 normal samples per category and diverse types of anomalies condensed into one anomaly class ‘bad’.
Evaluation Metrics
We assess the image-level detection and the pixel-level segmentation performance with three metrics each (similar to, e.g., Jeong et al., 2023; Chen et al., 2023; Li et al., 2024b). For evaluating the detection performance, we measure the area under the receiver-operator curve (AUROC), the F1-score at the optimal threshold (F1-max), and the average precision (AP) using the respective image-level anomaly scores. We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, Bergmann et al., 2021) of the segmentation using the pixel-wise anomaly scores. Note in this context that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% (Li et al., 2024b)—it is not recommended to assess performance solely on segmentation AUROC (Rafiei et al., 2023). We repeat each experiment three times and report mean and standard deviation. [^4] The randomness stems from the choice of reference samples $X_{\mathrm{ref}}$. For straightforward reproducibility, we simply set $X_{\mathrm{ref}}$ to the first $k$, second $k$ and third $k$ train samples for the three different runs, respectively.
[…]
Setting Method Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
1-shot SPADE 82.9$\pm$2.6 91.1$\pm$1.0 91.7$\pm$1.2 92.0$\pm$0.3 44.5$\pm$1.0 85.7$\pm$0.7
PatchCore 86.3$\pm$3.3 92.0$\pm$1.5 93.8$\pm$1.7 93.3$\pm$0.6 53.0$\pm$1.7 82.3$\pm$1.3
[…]
AnomalyDINO 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 96.3$\pm$0.1 58.2$\pm$0.8 91.9$\pm$0.1
2-shot SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
PatchCore 83.4$\pm$3.0 90.5$\pm$1.5 92.2$\pm$1.5 92.0$\pm$1.0 50.4$\pm$2.1 79.7$\pm$2.0
GraphCore 91.9$\pm$/ / / 96.9$\pm$/ / /
[…]
4-shot SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
PatchCore 88.8$\pm$2.6 92.6$\pm$1.6 94.5$\pm$1.5 94.3$\pm$0.5 55.0$\pm$1.9 84.3$\pm$1.6
GraphCore 92.9$\pm$/ / / 97.4$\pm$/ / /
[…]
8-shot GraphCore 95.9$\pm$/ / / 97.8$\pm$/ / /
WinCLIP+ 94.6$\pm$0.1 / / / / /
[…]
16-shot WinCLIP+ 94.8$\pm$0.1 / / / / /
[…]
AnomalyDINO 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.2$\pm$0.1 93.0$\pm$0.1
Baselines
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE (Cohen and Hoshen, 2020), PaDiM (Defard et al., 2020), PatchCode (Roth et al., 2022), WinCLIP+ (Jeong et al., 2023), and APRIL-GAN (Chen et al., 2023). It is important to note that ACR (Li et al., 2024a) and MuSc (Li et al., 2024b) consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported by Jeong et al. (2023) are indicated by , the result of the WinCLIP re-implementation by Kwak et al. (2024) (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP (Kwak et al., 2024) we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported (Xie et al., 2023).
4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively. Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ in every reported metric, outperforming even approaches that require training on an additional data set (e.g., ADP and APRIL-GAN).
[…]
Setting Method Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
1-shot SPADE 79.5$\pm$4.0 80.7$\pm$1.9 82.0$\pm$3.3 95.6$\pm$0.4 35.5$\pm$2.2 84.1$\pm$1.6
PaDiM 62.8$\pm$5.4 75.3$\pm$1.2 68.3$\pm$4.0 89.9$\pm$0.8 17.4$\pm$1.7 64.3$\pm$2.4
PatchCore 79.9$\pm$2.9 81.7$\pm$1.6 82.8$\pm$2.3 95.4$\pm$0.6 38.0$\pm$1.9 80.5$\pm$2.5
WinCLIP+ 83.8$\pm$4.0 83.1$\pm$1.7 85.1$\pm$4.0 96.4$\pm$0.4 41.3$\pm$2.3 85.1$\pm$2.1
APRIL-GAN 91.2$\pm$0.8 86.9$\pm$0.6 93.3$\pm$0.8 96.0$\pm$0.0 38.5$\pm$0.3 90.0$\pm$0.1
[…]
2-shot SPADE 80.7$\pm$5.0 81.7$\pm$2.5 82.3$\pm$4.3 96.2$\pm$0.4 40.5$\pm$3.7 85.7$\pm$1.1
PaDiM 67.4$\pm$5.1 75.7$\pm$1.8 71.6$\pm$3.8 92.0$\pm$0.7 21.1$\pm$2.4 70.1$\pm$2.6
PatchCore 81.6$\pm$4.0 82.5$\pm$1.8 84.8$\pm$3.2 96.1$\pm$0.5 41.0$\pm$3.9 82.6$\pm$2.3
[…]
ADP 86.9$\pm$0.9 / / / / /
APRIL-GAN 92.2$\pm$0.3 87.7$\pm$0.3 94.2$\pm$0.3 96.2$\pm$0.0 39.3$\pm$0.2 90.1$\pm$0.1
[…]
4-shot SPADE 81.7$\pm$3.4 82.1$\pm$2.1 83.4$\pm$2.7 96.6$\pm$0.3 43.6$\pm$3.6 87.3$\pm$0.8
PaDiM 72.8$\pm$2.9 78.0$\pm$1.2 75.6$\pm$2.2 93.2$\pm$0.5 24.6$\pm$1.8 72.6$\pm$1.9
[…]
ADP 88.4$\pm$0.4 / / / / /
APRIL-GAN 92.6$\pm$0.4 88.4$\pm$0.5 94.5$\pm$0.3 96.2$\pm$0.0 40.0$\pm$0.1 90.2$\pm$0.1
[…]
8-shot WinCLIP+ 85.0$\pm$0.0 / / / / /
ADP 89.2$\pm$0.1 / / / / /
[…]
16-shot WinCLIP+ 85.0$\pm$0.1 / / / / /
ADP 90.1$\pm$0.5 / / / / /
[…]
AnomalyDINO reduces the gap between few-shot and full-shot methods in anomaly detection significantly, and it demonstrates the strongest performance in anomaly localization. Furthermore, we see a clear improvement in the method’s results as the quantity of samples increases.
In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2,4\}$. Nonetheless, AnomalyDINO consistently achieves second-best results, comparable results with APRIL-GAN in the 8-shot setting, and sets new state-of-the-art for $k=16$. This can be attributed to AnomalyDINO’s ability to benefit more from a richer memory bank $\mathcal{M}$ than APRIL-GAN. We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples.Note also that, AnomalyDINO outperforms all other training-free approaches.
Table 4: Detection results for other settings. All results are AUROC values in (%).
Setting Method MVTec-AD VisA
0-shot WinCLIP 91.8 78.1
AnomalyCLIP 91.5 82.1
APRIL-GAN 86.1 78.0
Batched 0-shot ACR 85.8 /
MuSc 97.8 92.8
AnomalyDINO 93.3 89.6
Comparing the segmentation performance on VisA, Table 3 reveals a mixed picture: WinCLIP+ shows strong localization performance in terms of pixel-level AUROC, and APRIL-GAN achieves the highest scores in PRO. AnomalyDINO consistently achieves results in per-region-overlap that are second-best compared to all other, and the best among the training-free methods. In addition, AnomalyDINO shows the best segmentation performance when measured in F1-max, a metric that accounts for the strong imbalance of nominal and anomalous pixels.
We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D). This adaptation is straightforward and relatively simple, especially compared to MuSc. Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ (Li et al., 2024b). The results obtained, as detailed in Table 4, are therefore quite satisfactory.
We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings. Within the class of training-free models, it is the clear winner essentially across the board. These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures. We provide further qualitative results in Figures 3 and 4, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.
4.2 Ablation Study
We conduct additional experiments to assess the effect of specific design choices in our pipeline. The full results are provided in Appendix C. Here, we briefly describe the main insights.
Preprocessing
Comparing the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’ we observe that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 8 and 9. Depending on the product category suitable preprocessing steps lead to substantial improvements. In some cases augmenting the reference sample with rotations also negatively impacts the detection performance. For the full discussion, see Section C.1.
Inference time
Data augmentation increases the computational complexity of the proposed pipeline, as well as the memory bank size, and thus, the runtime of our nearest neighborhood approach. [^5] In practice, our implementation leverages GPU-accelerated neighborhood search (Johnson et al., 2019), leading to only slightly increased inference time. With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling (Roth et al., 2022) are advisable. The effects of masking and augmentations on the inference time, presented in Figure 11, are not negligible but can be reduced by approximately 30% in the ‘informed’ scenario (on MVTec-AD).
Aggregation statistic
We compare the default scoring method, the empirical tail value at risk, to other suitable aggregations in Section C.2. The mean of the 1% highest distances from test patches to $\mathcal{M}$ improves over the standard choice (maximum of the upsampled and smoothed patch distances).
Architecture size
In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 12. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k=16$). Interestingly, larger architectures do not necessarily translate to higher performance. As the dimensionality of the feature space $\mathcal{F}$ increases with larger architectures, the distances between nominal patches also increases. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture size L performs better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 (Oquab et al., 2024). We carefully design means to populate the nominal memory bank with diverse and relevant features utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
Follow-up research directions
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods (e.g., Fu et al., 2024) to obtain the anomaly map from patch features. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
Broader Impacts
Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle. Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring. It is, however, essential to recognize the shortcomings of automated anomaly detection systems. We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable. In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases for the proposed method). In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses. To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential. Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.
Appendix A Detailed experimental results
In this section, we present the results of the default setting (AnomalyDINO-S and ‘agnostic’ preprocessing) in detail. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6. Additional visualizations of anomaly maps are presented in Figures 3 and 4 for MVTec-AD and VisA, respectively.
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 3: Examples – MVTec-AD. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color coding is normalized by the maximum score on the ‘good’ test samples (per category).
Table 5: Detailed results on MVTec-AD. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
Shots 1-shot 2-shot 4-shot 8-shot 16-shot
AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
Bottle 99.5$\pm$0.4 96.1$\pm$0.3 99.7$\pm$0.1 96.4$\pm$0.1 99.9$\pm$0.1 96.6$\pm$0.2 100.0$\pm$0.0 96.6$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.2
Cable 91.6$\pm$1.2 88.2$\pm$0.7 92.1$\pm$0.7 88.5$\pm$0.7 93.8$\pm$1.4 89.5$\pm$0.5 94.7$\pm$0.3 89.6$\pm$0.3 94.4$\pm$1.1 89.5$\pm$0.7
Capsule 89.8$\pm$5.9 96.4$\pm$0.4 88.9$\pm$7.1 96.7$\pm$0.5 95.4$\pm$0.9 97.3$\pm$0.1 94.7$\pm$1.3 97.4$\pm$0.1 95.5$\pm$0.3 97.5$\pm$0.1
Carpet 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.2
Grid 99.2$\pm$0.1 96.1$\pm$0.0 99.2$\pm$0.1 96.3$\pm$0.1 99.2$\pm$0.1 96.2$\pm$0.2 99.4$\pm$0.2 96.2$\pm$0.1 99.8$\pm$0.3 95.9$\pm$0.3
Hazelnut 98.1$\pm$1.1 97.5$\pm$0.4 99.7$\pm$0.4 97.7$\pm$0.4 99.8$\pm$0.1 97.8$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.2 100.0$\pm$0.0 97.7$\pm$0.1
Leather 100.0$\pm$0.0 97.9$\pm$0.2 100.0$\pm$0.0 97.5$\pm$0.3 100.0$\pm$0.0 97.1$\pm$0.3 100.0$\pm$0.0 97.0$\pm$0.1 100.0$\pm$0.0 96.7$\pm$0.1
Metal nut 99.9$\pm$0.2 93.8$\pm$0.1 100.0$\pm$0.1 94.4$\pm$0.3 100.0$\pm$0.0 95.1$\pm$0.1 100.0$\pm$0.0 95.2$\pm$0.3 100.0$\pm$0.0 95.5$\pm$0.1
Pill 93.1$\pm$1.5 96.9$\pm$0.2 94.9$\pm$2.2 97.1$\pm$0.2 96.3$\pm$0.1 97.3$\pm$0.0 97.1$\pm$0.4 97.4$\pm$0.1 97.6$\pm$0.3 97.5$\pm$0.1
Screw 90.1$\pm$1.0 92.5$\pm$0.2 92.3$\pm$0.2 93.5$\pm$0.0 91.3$\pm$2.4 93.6$\pm$0.7 92.5$\pm$0.8 94.6$\pm$0.4 93.5$\pm$1.4 95.3$\pm$0.2
Tile 100.0$\pm$0.0 88.6$\pm$0.2 100.0$\pm$0.0 88.2$\pm$0.2 100.0$\pm$0.0 87.9$\pm$0.2 100.0$\pm$0.0 87.5$\pm$0.2 100.0$\pm$0.0 86.9$\pm$0.6
Toothbrush 99.4$\pm$0.5 94.5$\pm$0.1 99.9$\pm$0.1 94.9$\pm$0.3 99.9$\pm$0.1 95.0$\pm$0.3 99.9$\pm$0.1 95.5$\pm$0.5 99.6$\pm$0.3 95.9$\pm$0.6
Transistor 90.9$\pm$2.1 64.4$\pm$0.5 89.2$\pm$5.9 64.8$\pm$3.1 91.9$\pm$3.5 66.5$\pm$1.4 96.4$\pm$1.8 70.3$\pm$1.6 98.3$\pm$0.5 72.7$\pm$1.0
Wood 98.8$\pm$0.4 93.9$\pm$0.3 98.7$\pm$0.6 93.8$\pm$0.2 98.6$\pm$0.6 93.8$\pm$0.1 98.6$\pm$0.2 93.8$\pm$0.2 98.2$\pm$0.4 93.5$\pm$0.5
Zipper 98.9$\pm$0.1 83.2$\pm$0.4 99.0$\pm$0.1 84.4$\pm$0.7 99.3$\pm$0.2 85.6$\pm$0.2 99.2$\pm$0.2 85.5$\pm$0.6 99.3$\pm$0.3 85.9$\pm$0.7
Mean 96.6$\pm$0.5 91.9$\pm$0.1 96.9$\pm$0.8 92.1$\pm$0.2 97.7$\pm$0.1 92.5$\pm$0.1 98.2$\pm$0.1 92.8$\pm$0.0 98.4$\pm$0.1 93.0$\pm$0.1
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 4: Examples – VisA. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color bar is normalized by the maximum score on the ‘good’ test samples (per category). For the category ‘Capsules’ (left column, second from top) we changed the color map for better visibility. Best viewed at a higher zoom level as some anomalies are quite small.
Table 6: Detailed results on VisA. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
Shots 1-shot 2-shot 4-shot 8-shot 16-shot
AUROC PRO AUROC PRO AUROC PRO AUROC PRO AUROC PRO
Candle 83.8$\pm$0.7 77.6$\pm$0.6 86.2$\pm$3.8 78.4$\pm$0.1 88.0$\pm$3.0 79.3$\pm$0.3 89.9$\pm$1.0 79.7$\pm$0.2 92.0$\pm$0.8 80.7$\pm$0.3
Capsules 98.4$\pm$0.3 80.4$\pm$0.5 99.1$\pm$0.2 81.0$\pm$0.4 99.1$\pm$0.2 81.5$\pm$0.2 99.1$\pm$0.5 81.9$\pm$0.1 98.8$\pm$0.4 82.4$\pm$0.2
Cashew 85.3$\pm$5.6 95.0$\pm$1.5 89.3$\pm$5.6 95.7$\pm$1.1 93.8$\pm$0.5 96.2$\pm$0.8 95.1$\pm$0.7 95.9$\pm$0.3 95.5$\pm$0.1 96.2$\pm$0.3
Chewinggum 98.2$\pm$0.3 90.6$\pm$0.8 98.7$\pm$0.3 91.8$\pm$0.8 98.8$\pm$0.2 92.2$\pm$0.5 98.9$\pm$0.1 92.5$\pm$0.3 99.0$\pm$0.0 92.6$\pm$0.0
Fryum 95.1$\pm$0.3 92.6$\pm$0.3 96.6$\pm$0.5 93.2$\pm$0.1 97.2$\pm$0.5 93.6$\pm$0.1 97.9$\pm$0.6 93.8$\pm$0.1 97.8$\pm$0.6 94.0$\pm$0.1
Macaroni1 83.4$\pm$3.0 86.7$\pm$1.0 83.3$\pm$1.7 87.6$\pm$0.2 85.8$\pm$0.9 88.7$\pm$0.2 86.9$\pm$0.8 89.6$\pm$0.2 88.2$\pm$0.7 90.0$\pm$0.2
Macaroni2 61.4$\pm$3.8 66.3$\pm$0.7 67.1$\pm$5.3 67.5$\pm$0.3 70.0$\pm$1.8 68.5$\pm$0.4 74.0$\pm$1.2 70.3$\pm$0.6 76.5$\pm$1.5 71.9$\pm$0.5
PCB1 89.0$\pm$1.2 90.6$\pm$0.3 88.9$\pm$1.5 90.6$\pm$0.6 92.3$\pm$2.3 91.2$\pm$0.4 92.9$\pm$0.6 91.8$\pm$0.2 95.2$\pm$0.8 92.5$\pm$0.3
PCB2 87.8$\pm$2.0 86.5$\pm$1.1 89.4$\pm$2.5 87.5$\pm$0.2 91.8$\pm$0.7 88.4$\pm$0.1 92.8$\pm$0.6 89.1$\pm$0.2 93.0$\pm$0.4 89.5$\pm$0.3
PCB3 84.1$\pm$2.4 85.6$\pm$0.8 88.5$\pm$1.3 87.1$\pm$0.6 91.8$\pm$0.3 87.8$\pm$0.4 93.8$\pm$0.7 89.4$\pm$0.6 94.5$\pm$0.2 90.0$\pm$0.2
PCB4 76.7$\pm$11.5 79.8$\pm$5.0 86.5$\pm$7.2 83.3$\pm$3.7 94.9$\pm$3.0 86.1$\pm$0.8 97.3$\pm$0.6 88.8$\pm$1.1 99.0$\pm$0.3 90.6$\pm$0.9
Pipe fryum 91.8$\pm$2.6 97.4$\pm$0.3 93.5$\pm$3.0 97.5$\pm$0.3 94.8$\pm$2.7 97.2$\pm$0.1 96.5$\pm$1.6 97.3$\pm$0.1 98.0$\pm$0.3 97.3$\pm$0.1
Mean 86.2$\pm$1.4 85.8$\pm$0.6 88.9$\pm$1.9 86.8$\pm$0.4 91.5$\pm$0.7 87.6$\pm$0.2 92.9$\pm$0.1 88.3$\pm$0.2 94.0$\pm$0.0 89.0$\pm$0.2
The results presented in Tables 5 and 6 show that a) anomalies in some categories are more difficult to than others, b) detection and segmentation performance correlate strongly (measured in image-level AUROC and PRO), and c), that the performance of AnomalyDINO increases across the board with more available reference samples. In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples. In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s). The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
Appendix B Limitations and failure cases
The proposed method relies on similarities to patch representations captured in $\mathcal{M}$. Therefore, we can only expect the model to detect those anomalies caused by regions in the test images that are particularly different from patches of the given reference sample(s). Our experiments reveal that this may lead to some specific failure cases.
B.1 Semantic Anomalies
The first relates to the distinction between low-level sensory anomalies and high-level semantic anomalies. Semantic anomalies might be present because a logical rule or a specific semantic constraint is violated. In contrast, AnomalyDINO is developed with low-level sensory anomalies in mind.
(a) Reference sample
(b) Anomaly type ‘Bent Wire’
(c) Anomaly type ‘Cable Swap’
Figure 5: Example of a semantic anomaly in category ‘Cable’ (MVTec-AD). Depicted are a nominal reference sample, an example of a sensory anomaly (Figure 5(b)) and a semantic anomaly (Figure 5(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
Consider, for instance, the anomalies shown in Figure 5. While the cable with anomaly type ‘Bent Wire’ (Figure 5(b)) contains patches that cannot be well matched to patches of the reference image, this does not apply to the anomaly ‘Cable Swap’ (Figure 5(c)). The latter shows a semantic anomaly with two blue wires, while a nominal image of a ‘Cable’ should depict all three different cable types. As a result, all test patches of Figure 5(b) can be matched well with reference patches in $\mathcal{M}$, and thus, AnomalyDINO does not detect this anomaly type. Evaluating the detection performance for this specific failure case, ‘Cable Swap’, our proposed method essentially performs on chance level, giving a detection AUROC of 50.2% ($\pm$ 4.9%).
B.2 The importance of informative reference samples
The second failure case occurs if the reference sample(s) does not resemble all concepts of normality, therefore $\mathcal{M}$ does not capture all variations of the nominal distribution $p_{\mathrm{norm}}$. As an illustration, consider the nominal samples of ‘Capsule’ in MVTec-AD, which may be rotated in such a way that the imprinted text is hidden (see Figure 6(b)). Therefore, parts of the text of a nominal test sample may be falsely recognized as anomalies. Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 5 and 6. We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
(a) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
(b) Reference sample (left) and estimated anomaly map of ‘good’ test sample (right).
Figure 6: Example of an uninformative reference sample in category ‘Capsule’ (MVTec-AD). Some reference samples do not resemble the full concept of normality (here, the sample in Figure 6(b) does not show the text on the capsule such that any nominal sample with text will get a high anomaly score). Anomaly maps predicted by AnomalyDINO based on the depicted reference sample.
(a) Test sample (left), the same test sample with ground-truth anomaly annotation (center), and a magnified view of a region with strong background artifacts (right).
(b) Anomaly map predicted by AnomalyDINO (1-shot) without (left) and with masking (right).
Figure 7: Visualization of the effect of masking in the presence of high background noise for the category ‘Capsules’ in VisA. As in Figure 4 we depict the anomaly map for ‘Capsules’ using a different colormap (red instead of yellow) for better visibility. Best viewed on a higher zoom level.
Appendix C Ablation Study
C.1 Preprocessing
As discussed in Section 3, we consider two potential preprocessing steps in our pipeline: masking and rotations. We mask out irrelevant background patches, whenever the zero-shot segmentation of DINOv2 captures the object correctly (see Figure 2). Discarding background patches helps to mitigate the problem of potential background noise, thereby reducing the number of false positives. We find that some samples of anomalous samples also show artifacts in the background. An example of the category ‘Capsules’ from VisA is depicted in Figure 7. Without masking, the method would correctly predict the depicted sample to show an anomaly, but for the wrong reason (high anomaly scores for the background noise). We conclude that masking is essential to faithfully detect anomalies whenever higher variations in the background are expected. A quantitative analysis is provided in the next paragraphs.
In addition to masking (which decreases the size of $\mathcal{M}$), we consider augmenting the reference sample with rotations (which increases the size of $\mathcal{M}$). Here we distinguish the ‘agnostic’ scenario, where we do not know a priori about potential rotations, and augment by default. Whenever we know about potential rotations of reference or test samples, we can deactivate the augmentation to reduce the test time.
The preprocessing decision per category for MVTec-AD and VisA, inferred (solely) from the first nominal reference sample in the test set (to comply with the one-shot setting), are given in Table 7.
Table 7: Default preprocessing steps for MVTec-AD and VisA. We do not mask textures, indicated by (T). In addition, we do not apply masking when the masking test failed, indicated by (MT). See Section 3 and Figure 2 for further discussion and a visualization.
MVTec-AD Mask? Rotation?
Object informed agnostic
Bottle ✗(MT)
Cable ✗(MT)
Capsule
Carpet ✗(T)
Grid ✗(T)
Hazelnut
Leather ✗(T)
Metal nut ✗(MT)
Pill
Screw
Tile ✗(T)
Toothbrush
Transistor ✗(MT)
Wood ✗(T)
Zipper ✗(MT)
VisA Mask? Rotation?
Object informed agnostic
Candle
Capsules
Cashew
Chewinggum
Fryum
Macaroni1
Macaroni2
PCB1
PCB2
PCB3
PCB4
Pipe fryum
Effect of preprocessing on detection performance
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ influence the detection performance. The right preprocessing is in some cases more important than others. The full results are given in Figures 8 and 9.
Figure 8: Effect of preprocessing on MVTec-AD. Anomaly detection of AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
Figure 9: Effect of preprocessing on VisA. Anomaly detection with AnomalyDINO in the 1-shot setting for different choices of the preprocessing pipeline (detection AUROC on image-level in %, mean and standard deviation over three independent runs).
Rotation
Increasing the diversity of nominal patch representations in $\mathcal{M}$ by rotating the reference sample can significantly improve the detection performance. Consider, e.g., the category ‘Screw’ in MVTec-AD, where the detection AUROC can be boosted from 65.6% to 90.1% in the 1-shot setting. This is intuitive, as the test samples of ‘Screw’ are taken from various angles. With sufficiently many reference samples, such data augmentation is not necessary anymore, but for the few-shot setting, we see major improvements.
Figure 10: Rotations as anomalies (‘misplaced’ transistor from MVTec-AD). The left anomaly map is estimated from a reference sample with rotations (‘agnostic’), while the right anomaly map is based on a reference sample without rotations (‘informed’).
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, and ‘Metal nut’ (both MVTec-AD), as well as for ‘Macaroni2’, ‘Cashew’, ‘PCB2’ and ‘PCB4’ categories (all VisA). However, we also observe that rotations of the reference sample can actually decrease the detection performance in some categories, namely ‘PCB1’ and ‘PCB3’ (both VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 4, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$. Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 10. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 10).
Masking
Across the board, we see that masking improves the detection performance. The only exception is ‘Candle’ from VisA with a slight decrease in performance. On further inspection, we find that the estimated foreground masks of DINOv2 occasionally fail to capture all relevant patches (the decision on whether to mask or not is inferred from the first nominal sample only, 1-shot setting). We leave the development of more robust and adaptive masking techniques for future work.
Runtime analysis
The design choices in our pipeline influence the test time. To see the potential effects we measure the test time per sample on MVTec for the different combinations of masking (yes/no) and rotation (yes/no) for the full pipeline. This includes preparation of the memory bank $\mathcal{M}$ (i.e., feature extraction and the optional augmentation and masking operations), as well as the inference time for all test samples (i.e., feature extraction and optional masking). The results are given in Figure 11. The runtime was assessed utilizing a single NVIDIA A40 GPU, consistently employed throughout all experiments detailed in this paper (each experiment can be executed on a single NVIDIA A40 GPU).
The resulting run time increases by a factor of 4, when comparing the baseline (no masking, no rotation) to the scenario where masking and rotations are applied to each product. Not rotating a reference sample in cases where it is unnecessary (the difference between ‘informed’ and ‘agnostic’) lowers the test time by approximately 30%.
Figure 11: Effect on masking and augmentation on test time (per sample) in the 1-shot setting with AnomalyDINO. We measure the test time on a single NVIDIA A40 GPU and report mean and standard deviation over three independent runs on MVTec-AD.
C.2 Scoring
In Section 3 we investigate different ways of aggregating the anomaly scores $\mathcal{D}$ on patch-level (in our case, distances to the nominal memory bank $\mathcal{M}$) to an image-level anomaly score via a statistic $q$. Note that the segmentation results are therefore not affected by the choice of $q$.
A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow Roth et al., 2022, and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
Table 8: Effect of aggregation statistics $q$ on the detection performance on MVTec-AD and VisA. $\mathcal{D}$ denotes the set of patch distances to the nominal memory bank, $H_{0.01}(\mathcal{D})$ the 1% highest entries thereof, and $\mathcal{A}$ denotes the anomaly map derived from $\mathcal{D}$ (anomaly scores for each pixel, after upsampling and smoothing, see Section 3). All results in %.
Scoring $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ $q=\mathrm{max}(\mathcal{\mathcal{A}})$ $q=\mathrm{max}(\mathcal{D})$
Shots AUROC F1-max AP AUROC F1-max AP AUROC F1-max AP
MVTec-AD 1 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 95.8$\pm$0.6 95.4$\pm$0.3 97.7$\pm$0.4 95.1$\pm$0.5 94.8$\pm$0.1 97.4$\pm$0.4
2 96.9$\pm$0.8 96.6$\pm$0.4 98.2$\pm$0.7 96.2$\pm$1.1 96.0$\pm$0.4 97.7$\pm$0.7 95.6$\pm$1.0 95.4$\pm$0.5 97.4$\pm$0.6
4 97.7$\pm$0.1 97.1$\pm$0.3 98.5$\pm$0.3 97.2$\pm$0.2 96.7$\pm$0.3 98.1$\pm$0.4 96.7$\pm$0.2 96.1$\pm$0.1 97.9$\pm$0.4
8 98.2$\pm$0.1 97.5$\pm$0.1 99.1$\pm$0.2 97.7$\pm$0.1 97.1$\pm$0.3 98.8$\pm$0.1 97.3$\pm$0.2 96.5$\pm$0.4 98.6$\pm$0.1
16 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 98.0$\pm$0.1 97.4$\pm$0.2 99.1$\pm$0.1 97.7$\pm$0.2 96.9$\pm$0.3 98.9$\pm$0.1
VisA 1 86.2$\pm$1.4 83.5$\pm$1.0 87.7$\pm$1.4 85.1$\pm$1.7 82.7$\pm$1.0 87.1$\pm$1.3 83.8$\pm$1.8 82.0$\pm$1.0 85.8$\pm$1.6
2 88.9$\pm$1.9 85.3$\pm$1.4 90.4$\pm$1.4 87.7$\pm$1.9 84.3$\pm$1.4 89.7$\pm$1.2 86.6$\pm$1.9 83.4$\pm$1.2 88.5$\pm$1.2
4 91.5$\pm$0.7 87.6$\pm$0.9 92.6$\pm$0.6 90.7$\pm$0.3 86.4$\pm$0.6 92.1$\pm$0.3 89.8$\pm$0.3 85.7$\pm$0.5 91.2$\pm$0.5
8 92.9$\pm$0.1 89.0$\pm$0.3 93.8$\pm$0.1 92.0$\pm$0.1 87.8$\pm$0.3 93.2$\pm$0.3 91.2$\pm$0.1 87.0$\pm$0.2 92.4$\pm$0.4
16 94.0$\pm$0.0 90.0$\pm$0.1 94.8$\pm$0.1 93.3$\pm$0.2 88.9$\pm$0.4 94.4$\pm$0.2 92.6$\pm$0.3 88.2$\pm$0.4 93.7$\pm$0.3
The results in Table 8 show that just the maximum patch distance leads to already good results, but upsampling and smoothing the patch distances improves upon that (maximum of $\mathcal{A}$). Taking the mean of all patch distances above the 99% quantile ($q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$) improves above the standard choice. We did not optimize over the percentile and instead fixed it to 99%. Typically, the number of patches per image ranges between 200 and 1000 (depending on resolution and masking), such that 2 between 10 patches are considered. For specific products/objects and expected anomaly types, other statistics $q$ might be (more) suitable.
C.3 Architecture Size
DINOv2 is available in different distillation sizes (S, B, L, and G). This section analyzes the implications of choosing different backbone sizes for AnomalyDINO. The comparison including the best competing methods is depicted in Figure 12.
We see that different architecture sizes have some effect on the image-level AUROC. AnomalyDINO-L outperforms smaller architectures, S (our default) and B, on VisA in the few-shot setting with $k\in\{1,2,4\}$. For $k\in{8,16}$ AnomalyDINO-L is on par with AnomalyDINO-S (and APRIL-GAN). For $k=16$ all architecture sizes outperform the best competitor (APRIL-GAN).
On MVTec-AD, we see that AnomalyDINO-S performs slightly better, but the gap between AnomalyDINO-S and AnomalyDINO-B reduces with increasing shot size. In particular, all architecture sizes outperform the closest competitor (ADP).
Refer to captionRefer to caption
Figure 12: Effect of model size on detection AUROC. We report mean and standard deviation for the 1-shot setting on MVTec-AD (left) and VisA (right) over three seeds. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by Kwak et al. (2024).
Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We extend the proposed method to the batched zero-shot setting. Recall, that in this setting all test samples $X_{\mathrm{test}}$ are provided (or at least a sufficiently large batch), but no (labeled) training or reference samples. The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., Li et al., 2024a).
We need to alter our method, outlined in Section 3, only slightly to adapt it to the batched setting. We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring (Li et al., 2024b). For each test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ we therefore collect all patch representations not belonging to $\mathbf{x}^{(j)}$ in a memory bank, again utilizing DINOv2 as patch-level feature extractor $f$,
$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in\{1,\dots,n\}\bigr{\}}\enspace.$ (5)
We need to infer anomaly scores for each patch representation $\mathbf{p}_{\mathrm{test}}$ of $\mathbf{x}^{(j)}$. We could again assess the distances between $\mathbf{p}_{\mathrm{test}}$ and $\mathcal{M}_{j}$ based on the distance to the nearest neighbor, as done in Equation 2. Note, however, that $\mathcal{M}_{j}$ may now contain nominal and anomalous patches. Thus, the nearest-neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.
A simple solution, based on the assumption that the majority of patches are nominal, is to replace the nearest neighbor with a suitable aggregation statistic over the distribution of patches distances in $\mathcal{M}_{j}$
$\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j}):=\big{\{}d(\mathbf{p}_{\mathrm{test}},\mathbf{p})\mid\mathbf{p}\in\mathcal{M}_{j}\big{\}}\enspace,$ (6)
where we again use the cosine distance $d$, defined in Equation 3. Specifically, we can again make use of the tail value at risk—now for the lowest quantile as we are interested in the tail behavior of the distribution of distances to the nearest neighbors—to derive a patch-level anomaly score
$s(\mathbf{p}_{\mathrm{test}}):=\mathrm{mean}\big{(}L_{\alpha}\big{(}\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})\big{)}\big{)}\enspace,$ (7)
where $L_{\alpha}(\mathcal{D})$ contains the values below the $\alpha$ quantile of $\mathcal{D}$. We set $\alpha=0.1\%$ as anomalous patches are rare by assumption, and $\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})$ large enough to accurately estimate the tail of the distribution. [^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464. The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is again given by aggregating the patch-level anomaly score following Equation 4. Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 0.1 seconds/sample for MVTec-AD (again measured on an NVIDIA A40; feature extraction takes approximately the same amount of time).
Some resulting anomaly maps in the batched zero-shot setting are depicted in Figure 13.
Figure 13: Anomaly maps for the batched zero-shot setting on MVTec-AD. The left-most sample in each category is a ‘good’ test sample for reference, followed by four samples with anomalies.